OpenStudy (nevermind_justschool):

Dylan opened a credit card account with $650.00 of available credit. Now that he has made some purchases, Dylan's account only has $377.00 of available credit. What is the percentage decrease of the amount of available credit in Dylan's account? 58% 32% 72% 42%

OpenStudy (texaschic101):

(650 - 377) / 650 = 273/650 = 0.42 = 42%

OpenStudy (nevermind_justschool):

@bohotness

OpenStudy (bohotness):

O-O WHNAT IS THAT?

OpenStudy (bohotness):

IS IT IN MY CLASS OR THE 7TH GRADE THING?????????????

OpenStudy (nevermind_justschool):

THIS LOOK AT IT

OpenStudy (nevermind_justschool):

Did you do it yet????

OpenStudy (bohotness):

I DONT SEE IT O-O ITS NOT ON MY THING O-O I AM OVCA

OpenStudy (nevermind_justschool):

Good, I guess I'm in a different thingy phew I'm in GCA

OpenStudy (bohotness):

http://prntscr.com/6lyz1s <------- HELP MEEEEEEEEEEEEEEEEE

OpenStudy (nevermind_justschool):

Sorry, but I haven't even started my math posttest :(

OpenStudy (bohotness):

okay i iwl lhelp you CX

OpenStudy (nevermind_justschool):

Really??